OpenStudy (anonymous):

What property is this equation: 10(xy+v)=10xy+10v

OpenStudy (anonymous):

options pls (:

OpenStudy (anonymous):

a) commuatative property of multiplication b)associative property of addition c)associative property of multiplication d)distributive property

OpenStudy (sburchette):

It is the distributive property. In general terms it states\[a(b+c)=ab+ac\]

OpenStudy (anonymous):

so would 7(8x)=(7*8)x be associative property of multiplication

OpenStudy (sburchette):

Yes, that's associative.

OpenStudy (anonymous):

thank yo